OpenStudy (anonymous):

What is the 10th term of the sequence 64, 16, 4, ….?

hartnn (hartnn):

do you find any pattern in this sequence ?

OpenStudy (wolf1728):

WeNeedToTalk 64 is 4³ 16 is 4² See a pattern now?

OpenStudy (anonymous):

I see a pattern but I'm not sure what it means. So would I do 4^-10 ?

OpenStudy (wolf1728):

You wouldn't have to calculate 4^-10 Assuming 64 is the first number in the sequence then you would carry it out to 4^-6 I guess you don't know negative exponents. 4^-2 is 1/4² = 1/16

hartnn (hartnn):

If this question belongs to Geometric series topic, then we can also start by finding the common ratio 'r' common ratio is the ratio between the terms, r = 2nd term/1st term = 3rd term/2nd term = ... and so on here, r = 16/64 = 4/16 = 1/4 now just use the formula for n'th term, \(\Large a_n = a_1 r^{n-1}\) where a1 = 1st term = 64 r= 1/4 n=10 plug in! :)

OpenStudy (anonymous):

@hartnn yea thats what I'm learning thanks!

hartnn (hartnn):

then this might be helpful to you :) http://openstudy.com/study#/updates/503bb2a0e4b007f9003103b0

OpenStudy (wolf1728):

Gee - guess I was concentrating too much on the numeric aspects of this. :-)